OpenStudy (maddytarin_98):

The square of 9 less than a number is 3 less than the number. What is the number? –12 or 7 –12 or –7 –7 or 12 7 or 12

OpenStudy (maddytarin_98):

i appreciate interaction and the help but answers are greatly appreciated. Imtaking a timed quiz,

OpenStudy (slenderman):

so 3 less than a number is 3 less than a number, well that at least mewans its positive.

OpenStudy (slenderman):

or is it a number minus 9^2 is 3 less than a number either ways it is still gonna be positive

OpenStudy (yanasidlinskiy):

(x - 9)² = x - 3 x² - 18x + 81 = x - 3 x² - 19x + 84 = 0 (x - 12)(x - 7) = 0 x = 12 or x = 7

OpenStudy (yanasidlinskiy):

Right. They both will be postitive:)

OpenStudy (maddytarin_98):

thank you guys! @Slenderman @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Yeap